What’s the current yield of a 5.70 percent coupon corporate bond quoted at a price of 97.28?
Answer #1

Annual coupon=(100*5.7%)=$5.7

Hence current yield=Annual coupon/Current price

=(5.7/97.28)

which is equal to

=5.86%(Approx).